Setting Standards, Materials and Labor Variances

Tom Belford and Tony Sorrentino own a small business devoted to kitchen and bath granite installations. Recently, building contractors have insisted on up-front bid prices for a house rather than the cost-plus system that Tom and Tony were used to. They worry because natural flaws in the granite make it impossible to tell in advance exactly how much granite will be used on a particular job. In addition, granite can be easily broken, meaning that Tom or Tony could ruin a slab and would need to start over with a new one. Sometimes the improperly cut pieces could be used for smaller installations, sometimes not. All their accounting is done by a local certified public accounting firm headed by Charlene Davenport. Charlene listened to their concerns and suggested that it might be time to implement tighter controls by setting up a standard costing system.

Charlene reviewed the invoices pertaining to a number of Tom and Tony's previous jobs to determine the average amount of granite and glue needed per square foot. She then updated prices on both materials to reflect current conditions. The standards she developed for one square foot of counter installed were as follows:

Granite, per square foot $50.00
Glue (10 oz. @ $0.15) 1.50
Direct labor hours:
Cutting labor (0.10 hr. @ $15) 1.50
Installation labor (0.25 hr. @ $25) 6.25

These standards assumed that one seamless counter requires one sink cut (the space into which the sink will fit) as well as cutting the counter to fit the space available.

Charlene tracked the actual costs incurred by Tom and Tony for granite installation for the next six months. She found that they completed 50 jobs with an average of 32 square feet of granite installed in each one. The following information on actual amounts used and cost was gathered:

Granite purchased and used (1,640 sq. ft.) $79,048
Glue purchased and used (16,000 oz.) $ 2,560
Actual hours cutting labor 180
Actual hours installation labor 390

The actual wage rate for cutting and installation labor remained unchanged from the standard rate.

Required:

If an amount is zero, enter "0" and select neither from the dropdown.

1. Calculate the materials price variances and materials usage variances for granite and for glue for the past six months.

Granite
Materials price variance $
Materials usage variance $
Glue
Materials price variance $
Materials usage variance $

2. Calculate the labor rate variances and labor efficiency variances for cutting labor and for installation labor for the past six months.

Cutting Labor
Labor rate variance $
Labor efficiency variance $
Installation Labor
Labor rate variance $
Labor efficiency variance $

3. Conceptual Connection: Would it be worthwhile for Charlene to establish standards for atypical jobs (such as those with more than one sink cut or wider than normal)?
Answer #1

1)

Granite
Materials price variance =79,048 - (1640 x 50) 2,952.00 F
Materials usage variance=(1640 - 1600)50 2,000.00 U
Glue
Materials price variance = 2560 - (16000 x 0.15)     160.00 U
Materials usage variance (16000 - 16000) 015           -  

2)

Cutting Labor
Labor rate variance = (15 - 15)180         -  
Labor efficiency variance = (180 - 160)15 300.00 U
Installation Labor
Labor rate variance = (25 - 25)390         -  
Labor efficiency variance = (390 - 400)25 250.00 f

3) No it Would not be worthwhile for Charlene to establish standards for a typical jobs , the operation is small which can be decided mentally also